OpenStudy (anonymous):

32z + w + t = 66 when w = 1 and t = 1

OpenStudy (anonymous):

what's z=

OpenStudy (anonymous):

substitude 1 for w and t 32z + 1 + 1 = 66 subtract 2 from both sides 32x + 2 -2 = 66 -2 32x = 64 divide by sides by 32 32x/32 = 64/32 x = 2

OpenStudy (anonymous):

i mean z = 2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thx so i dontget it why do they only give me 2 numbers for 3 variables??

OpenStudy (anonymous):

so you can practice your algebra you need at least one equation for every variable to solve a system of equations. they give you two so you only have to solve for one variable.

OpenStudy (anonymous):

ok